# A novel multipoint stress control volume method for linear elasticity on quadrilateral grids

### v1

@@ 5.1 Darcy flow
where $A$ is the mass matrix for $\bm{u}_{h}$ corresponding to each interaction region and $p=-(p_{2}-p_{1},p_{3}-p_{2},p_{3}-p_{4},p_{4}-p_{1})$; see Figure 3. We can observe that this elimination is in a similar fashion to that of [26, 3.13].
@@ 6 Numerical experiments
In this section, we present several examples to verify the convergence of our method. Specifically, the relative errors of displacement $\bm{u}_{h}$, stress tensor $\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ and the corresponding convergence rates will be reported. For simplicity of notation, we denote $\|a-b\|_{L^{2}(\Omega)}$ as the relative $L^{2}$-error between $a$ and $b$ instead of using $\|a-b\|_{L^{2}(\Omega)}/\|b\|_{L^{2}(\Omega)}$. Moreover, $\underline{\sigma}_{h}$ has different values in subcells of an macro-element, we compute the mean value of $\underline{\sigma}_{h}$ over a macro-element denoted by $\mathcal{M}\underline{\sigma}_{h}$ and its relative error against exact solution in the center of each macro-element. In the first example, the analytical solution is given by
@@
with Dirichlet boundary condition, and the body force is then computed using Lamé coefficients $\lambda=123,\mu=79.3$. Table 1 shows errors and convergence rates on a sequence of uniform mesh refinements, computed using rotation over the interaction region (Method 1) and macro-element (Method 2). Convergence rates of $\bm{u}_{h}$ and $\underline{\sigma}_{h}$ agree with the theoretical error analysis. Superconvergence of $\mathcal{M}\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ are observed. Figure 4 shows an example of residual distribution which clearly demonstrates the local conservation property (see Remark 3.1) of the proposed method. The second example aims to illustrate the convergence behavior of our methods in three dimension, in particular, the analytical solution is
@@
We set $\kappa=10^{6}$ and $\lambda=\mu=(1-\chi)+\kappa\chi$. Since the coefficient here is highly heterogeneous, we consider the modified weak formulation (3.5)-(3.7) for Method 1. Table 3 lists errors and convergences rates of this heterogeneous example. Similar convergence rates as example 2 can be found, it is worthwhile to mention superconvergence of $\mathcal{M}\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ are also observed. The computed displacement for this example are displayed in the left panel of Figure 5. Example 4 intends to check the locking-free property of the proposed method, specifically, the exact solution is given by
@@
which yields a deformed computational domain with $4\times 4$ quadrilateral grid. We then apply a uniform refinement of this $4\times 4$ quadrilateral grid to generate a sequence of grids (see middle panel of Figure 5), the resulting sequence of mesh satisfies the $h^{2}$-parallelogram property. Convergence rates on this grid are shown in Table 5, at least first-order convergence for all variables in their respective norms are observed, there are no significant degeneration of the convergence rate. For the second unstructured grid test, we compute solution on a sequence of smooth quadrilateral meshes which is obtained by applying a smooth map
@@
to a uniformly refined square mesh (see right panel of Figure 5). We reported errors and convergence rates on this meshes in Table 6, similar convergence behaviors as parallelogram mesh are found. Displacement computed on this type of mesh are depicted in right panel of Figure 5.
We also test our method on sequences of quadrilateral grids generated by random perturbation of uniform grids. Specifically, at each refinement level, the vertices are randomly perturbed within a circle whose radius is $h^{2}$. Results for these meshes are reported in Table 7, again, the proposed method delivers superconvergence of displacement and at least first-order convergence of other variables.
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 2.1107E-01 / 9.2463E-02 / 1.3043E-01 / 1.5986E-01 / $1/8$ 1.0759E-01 0.9722 3.2100E-02 1.5263 3.1790E-02 2.0366 6.3990E-02 1.3209 $1/16$ 5.3408E-02 1.0104 8.8784E-03 1.8542 8.0459E-03 1.9822 2.7066E-02 1.2414 $1/32$ 2.6938E-02 0.9874 2.5908E-03 1.7769 2.0474E-03 1.9745 9.9050E-03 1.4503 $1/64$ 1.3501E-02 0.9966 7.8447E-04 1.7236 5.1790E-04 1.9830 3.3807E-03 1.5508 $1/128$ 6.7604E-03 0.9979 2.4824E-04 1.6600 1.3003E-04 1.9938 1.1457E-03 1.5611 $1/256$ 3.3818E-03 0.9993 8.1945E-05 1.5990 3.2551E-05 1.9981 3.9327E-04 1.5426 $1/512$ 1.6913E-03 0.9997 2.7889E-05 1.5550 8.1419E-06 1.9993 1.3672E-04 1.5243
@@
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 2.2844E-01 / 1.2264E-01 / 1.4740E-01 / 2.4499E-01 / $1/8$ 1.3232E-01 0.7878 4.8392E-02 1.3416 4.7768E-02 1.6256 1.3381E-01 0.8725 $1/16$ 6.7906E-02 0.9624 1.6935E-02 1.5148 1.4576E-02 1.7124 6.6473E-02 1.0093 $1/32$ 3.4512E-02 0.9764 5.3464E-03 1.6634 4.0691E-03 1.8408 2.3594E-02 1.4943 $1/64$ 1.7404E-02 0.9877 1.6989E-03 1.6540 1.0603E-03 1.9402 7.5525E-03 1.6434 $1/128$ 8.7378E-03 0.9941 5.6783E-04 1.5811 2.6858E-04 1.9810 2.4374E-03 1.6316 $1/256$ 4.3773E-03 0.9972 1.9464E-04 1.5447 6.7426E-05 1.9940 8.1185E-04 1.5861 $1/512$ 2.1907E-03 0.9986 6.7785E-05 1.5218 1.6881E-05 1.9979 2.7740E-04 1.5492
@@
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 1.9552E-01 / 9.8976E-02 / 1.3217E-01 / 2.1107E-01 / $1/8$ 1.0162E-01 0.9441 2.7386E-02 1.8536 2.8569E-02 2.2099 4.5648E-02 2.2091 $1/16$ 4.9652E-02 1.0333 7.5703E-03 1.8550 7.1306E-03 2.0024 1.3117E-02 1.7991 $1/32$ 2.4909E-02 0.9952 2.1222E-03 1.8348 1.7565E-03 2.0213 3.6408E-03 1.8491 $1/64$ 1.2447E-02 1.0009 6.7770E-04 1.6468 4.3809E-04 2.0034 1.2704E-03 1.5190 $1/128$ 6.2262E-03 0.9994 2.8947E-04 1.2272 1.0955E-04 1.9996 5.3802E-04 1.2396 $1/256$ 3.1137E-03 0.9997 1.3555E-04 1.0946 2.7394E-05 1.9997 2.5804E-04 1.0601 $1/512$ 1.5568E-03 1.0000 6.6942E-05 1.0178 6.8473E-06 2.0003 1.2701E-04 1.0227

### v2

@@ 5.1 Darcy flow
where $A$ is the mass matrix for $\bm{u}_{h}$ corresponding to each interaction region and $p=-(p_{2}-p_{1},p_{3}-p_{2},p_{3}-p_{4},p_{4}-p_{1})$; see Figure 3. We can observe that this elimination is in a similar fashion to that of [23, 3.13].
@@ 6 Numerical experiments
In this section, we present several examples to verify the convergence of our method. Specifically, the relative errors of displacement $\bm{u}_{h}$, stress tensor $\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ and the corresponding convergence rates will be reported. For simplicity of notation, we denote $\|a-b\|_{L^{2}(\Omega)}$ as the relative $L^{2}$-error between $a$ and $b$ instead of using $\|a-b\|_{L^{2}(\Omega)}/\|b\|_{L^{2}(\Omega)}$. Moreover, $\underline{\sigma}_{h}$ has different values in subcells of an macro-element, we compute the mean value of $\underline{\sigma}_{h}$ over a macro-element denoted by $\mathcal{M}\underline{\sigma}_{h}$ and its relative error against exact solution in the center of each macro-element.
6.1 Example 1: 2D structured grid case
In the first example, the analytical solution is given by
@@ 6.1 Example 1: 2D structured grid case
with Dirichlet boundary condition, and the body force is then computed using Lamé coefficients $\lambda=123,\mu=79.3$. Table 1 shows errors and convergence rates on a sequence of uniform mesh refinements, computed using rotation over the interaction region (Method 1) and macro-element (Method 2). Convergence rates of $\bm{u}_{h}$ and $\underline{\sigma}_{h}$ agree with the theoretical error analysis. Superconvergence of $\mathcal{M}\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ are observed. Figure 4 shows an example of residual distribution which clearly demonstrates the local conservation property (see Remark 3.1) of the proposed method.
6.2 Example 2: 3D structured grid case
The second example aims to illustrate the convergence behavior of our methods in three dimension, in particular, the analytical solution is
@@ 6.3 Example 3: heterogeneous materials
We set $\kappa=10^{6}$ and $\lambda=\mu=(1-\chi)+\kappa\chi$. Since the coefficient here is highly heterogeneous, we consider the modified weak formulation (3.5)-(3.7) for Method 1. Table 3 lists errors and convergences rates of this heterogeneous example. Similar convergence rates as example 2 can be found, it is worthwhile to mention superconvergence of $\mathcal{M}\underline{\sigma}_{h}$ and rotation $\gamma_{h}$ are also observed. The computed displacement for this example are displayed in the left panel of Figure 6.
6.4 Example 4: nearly incompressible materials
Example 4 intends to check the locking-free property of the proposed method, specifically, the exact solution is given by
@@ 6.5 Example 5: 2D unstructured grid cases
which yields a deformed computational domain with $4\times 4$ quadrilateral grid. We then apply a uniform refinement of this $4\times 4$ quadrilateral grid to generate a sequence of grids (see an example in Figure 5(a)), the resulting sequence of mesh satisfies the $h^{2}$-parallelogram property. Convergence rates on this grid are shown in Table 5, at least first-order convergence for all variables in their respective norms are observed, there are no significant degeneration of the convergence rate. For the second unstructured grid test, we compute solution on a sequence of smooth quadrilateral meshes which is obtained by applying a smooth map
@@
to a uniformly refined square mesh (see an example in Figure 5(b)). We reported errors and convergence rates on this meshes in Table 6, similar convergence behaviors as parallelogram mesh are found. Displacement computed on this type of mesh are depicted in right panel of Figure 6.
We also test our method on sequences of quadrilateral grids generated by random perturbation of uniform grids (see an example in Figure 5(c)). Specifically, at each refinement level, the vertices are randomly perturbed within a circle whose radius is $h^{2}$. Results for these meshes are reported in Table 7, again, the proposed method delivers superconvergence of displacement and at least first-order convergence of other variables.
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 2.1107E-01 / 9.2463E-02 / 1.3043E-01 / 1.5986E-01 / $1/8$ 1.0759E-01 0.9722 3.2100E-02 1.5263 3.1790E-02 2.0366 6.3990E-02 1.3209 $1/16$ 5.3408E-02 1.0104 8.8784E-03 1.8542 8.0459E-03 1.9822 2.7066E-02 1.2414 $1/32$ 2.6938E-02 0.9874 2.5908E-03 1.7769 2.0474E-03 1.9745 9.9050E-03 1.4503 $1/64$ 1.3501E-02 0.9966 7.8447E-04 1.7236 5.1790E-04 1.9830 3.3807E-03 1.5508 $1/128$ 6.7604E-03 0.9979 2.4824E-04 1.6600 1.3003E-04 1.9938 1.1457E-03 1.5611
@@
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 2.2844E-01 / 1.2264E-01 / 1.4740E-01 / 2.4499E-01 / $1/8$ 1.3232E-01 0.7878 4.8392E-02 1.3416 4.7768E-02 1.6256 1.3381E-01 0.8725 $1/16$ 6.7906E-02 0.9624 1.6935E-02 1.5148 1.4576E-02 1.7124 6.6473E-02 1.0093 $1/32$ 3.4512E-02 0.9764 5.3464E-03 1.6634 4.0691E-03 1.8408 2.3594E-02 1.4943 $1/64$ 1.7404E-02 0.9877 1.6989E-03 1.6540 1.0603E-03 1.9402 7.5525E-03 1.6434 $1/128$ 8.7378E-03 0.9941 5.6783E-04 1.5811 2.6858E-04 1.9810 2.4374E-03 1.6316
@@
(a) Parallelogram mesh, $h=1/32$.
(b) Smooth quadrilateral mesh, $h=1/32$.
(c) Randomly perturbed grids $h^{2}$, $h=1/32$.
Figure 5: Examples of unstructured meshes.
@@
$h$ $\|\underline{\sigma}-\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\underline{\sigma}-\mathcal{M}\underline{\sigma}_{h}\|_{L^{2}(\Omega)}$ $\|\bm{u}-\bm{u}_{h}\|_{L^{2}(\Omega)}$ $\|\gamma-\gamma_{h}\|_{L^{2}(\Omega)}$ Error Rate Error Rate Error Rate Error Rate $1/4$ 1.9552E-01 / 9.8976E-02 / 1.3217E-01 / 2.1107E-01 / $1/8$ 1.0162E-01 0.9441 2.7386E-02 1.8536 2.8569E-02 2.2099 4.5648E-02 2.2091 $1/16$ 4.9652E-02 1.0333 7.5703E-03 1.8550 7.1306E-03 2.0024 1.3117E-02 1.7991 $1/32$ 2.4909E-02 0.9952 2.1222E-03 1.8348 1.7565E-03 2.0213 3.6408E-03 1.8491 $1/64$ 1.2447E-02 1.0009 6.7770E-04 1.6468 4.3809E-04 2.0034 1.2704E-03 1.5190 $1/128$ 6.2262E-03 0.9994 2.8947E-04 1.2272 1.0955E-04 1.9996 5.3802E-04 1.2396